there are 155 coin nickels,dimes, quarters= 24.95 the number of dimes is t more than twice the number of nickels
Answer by ankor@dixie-net.com(22740)

Write an equation for each phrase:
:
"there are 155 coin nickels,dimes, quarters"
n + d + q = 155
:
"nickels, dimes, quarters = 24.95 "
.05n + .10d + .25q = 24.95
:
"the number of dimes is two more than twice the number of nickels"
d = 2n + 2
:
Replace (2n+2) for d in the coin# and $amt equations:
:
1st equation:
n + (2n+2) + q = 155
3n + 2 + q = 155
3n + q = 155 - 2
3n + q = 153
:
2nd Equation
.05n + .10(2n+2) + .25q = 24.95
.05n + .2n + .2 + .25q = 24.95
.25n + .25q = 24.95 - .2
.25n + .25q = 24.75
:
Multiply the 2nd equation by 4 and subtract from the 1st equation
3n + q = 153
1n + q = 99
-------------subtracting eliminates q
2n = 54
n = 54/2
n = 27 nickels
:
Given d = 2n+2
d = 2(27) + 2
d = 56 dimes
:
Find q using the coin# equation
27 + 56 + q = 155
q = 155 - 83
q = 72 quarters
:
:
Check solution using $amt equation:
.05(27) + .10(56) + .25(72) = 24.95
:
Did this help you understand this?